A chain of length
L and mass density ¾
is held such that it hangs vertically just
above a scale. It is then released. What is the reading on the scale, as a function of
the height of the top of the chain?

Grade:

2 Answers

Arun Kumar IIT Delhi
9 years ago
Hello
Well most probably it will depend linearly with respect to length of the chain.
But a graphic explanation of the problem could have been led to a specific solution.
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Kaviraj
6 years ago
Well , the question is one with variable mass, so either you can write the equation of F =dp/dt or solve it in parts,(k is the density)
SOLUTION:
 
Let height be x at a certain instant ,
Writing the equation,
F-kgL  = kax+kv2
here a will be equal to g as it is freefall and v will the velocity that comes from energy conservation of the top part of the chain.
Solving we get F = 3kg(L-x) =9kg(L-x)/4